# Zero Trust Architecture: A Systematic Literature Review

### v1

@@ 5.1.10 Smart Energy Systems
Smart Energy Systems (SES) integrate modern technologies like the Internet of Things (IoT) [154], cloud computing, and artificial intelligence into existing energy infrastructure. Such systems are meant to maximize energy generation, distribution, and consumption while improving reliability and sustainability of power [155]. However, this interconnectedness brings significant cyber security risks, adversaries could take advantage of vulnerabilities to disrupt services or create economic and societal damage [156]. In essence, ZTA can serve as a powerful model to tackle these security challenges as it guarantees that only authenticated and authorized parties such as devices, users, or systems can access sensitive energy infrastructure [26]. Out of the reviewed articles, three specific areas have been identified in which ZTA enhances the security of intelligent energy systems.
@@
Furthermore, Hrishav et al. put forward a ZTA framework based on ABAC to bolster the security of AMI within smart grids [161]. The framework employs a trust score-based methodology, where each attribute (a set of 15 compiled attributes) receives a trust score which signifies the level of trust put in it and a corresponding threshold score which is represents the minimum required trust score for a given attribute to be taken as reliable. The authors consider a case study using five attributes which include ”Smart Meter ID, Device Version, Location, Challenge Response Pairs (CRPs)”, and ”Network Used” where a smart meter starts an access request to transmit usage data to a power supply company. Network access is granted only when the total trust score value is more than the total threshold value of the attributes. The case study shows the importance of each attribute in the access verification process within the AMI system. Also, the verification mechanism, driven by trust scores, ensures secure and reliable communication between smart meters and the utility company which effectively prevent unauthorized access and maintaining data integrity.
@@
Power IoT devices such as sensors and controllers are used in smart grids for optimal energy management, however, these devices are also vulnerable to cyber threats like malware and denial of service attack. Through micro-segmentation and other strategies, ZTA can protect power IoT systems such that a compromised device cannot propagate its effect to the full network [119]. Based on the key principles of ZT and the features of the power IoT, the authors in [154] proposed the ZT framework that can evaluate the security state of power IoT derived from the behaviuor and characteristics of the terminals. Through a comparative study based on simulation results and stochastic Petri nets, they prove that the proposed framework can be used to effectively solve network security problems, and significantly improve the security protection of the power IoT. Similarly, to fortify the access control and protect power IoT environment against and disgruntle users and compromised devices, a ZTA framework based on attribute-based encryption method is proposed in [78]. It is made up of three interconnected modules - ”control plane” acting as the intelligent core of the framework and responsible for making real-time access decisions through dynamic trust evaluation and security access control components, ”data plane” which receives and executes the access control decisions from the control plane, and the ”identity security infrastructure” which handles the registration and authentication process of users and devices, manages their attributes, and defines the access policies that govern access to specific resources. The interaction of these three modules ensures that only trusted entities with validated identities and acceptable trust scores are granted access to sensitive data within the power IoT environment. Experimental results show that our framework is effective against the identified security vulnerabilities.
@@
Development in information technology (IT) has both significantly improve the power grids and systems operations and exposed the power infrastructure to diverse security challenges which the traditional security controls cannot be able to cope with effectively [162]. Hence, robust and dynamic security solutions such as ZTA frameworks are needed to secure data and information and terminal devices in the power system networks [163]. For example, in [163], the authors proposed a three-pronged approach to developing a ZTA framework for securing terminal access in distributed energy grids: ”Security modules”, like ”Trusted Platform Modules (TPMs) or Hardware Security Modules (HSMs)”, act as trust anchors for the system storing important information like digital certificates or terminal keys. ”Identity authentication” which relies on lightweight one-time digital signatures, ensuring real-time authentication and enhancing security by generating a new signature for each request. ”Security detection” employs data stream classification to continuously monitor the security of terminals, facilitating real-time analysis and anomaly detection. Through comparative analysis, the proposed framework is shown to outperformed traditional methods and can address the security challenges of decentralized power grids.
@@ 5.2 Application Requirements
In order to successfully deploy ZTA, specific baseline requirements must be met. Micro-segmentation is a cornerstone, dividing networks into smaller, isolated zones to minimize the impact of breaches and prevent lateral movement of threats [14], [106]. This includes specifying trust dimensions for each access request such that access is granted only when defined security parameters are satisfied [21]. Continuous monitoring and verification is essential to the ZTA model by analyzing user and device activities in real time to detect anomalies and enforce security policies dynamically [7]. This process relies on monitoring and analytics solutions, which involve the use of Security Information and Event Management (SIEM) [73] and Security Orchestration, Automation and Response (SOAR) systems [164] to provide centralized log data aggregation from systems and appliances, real-time threat detection, and automated response [165]. Security concepts such as threat intelligence platforms and activity log management further enhance situational awareness by providing insights into emerging threats and enabling traceability for incident response [6]. This includes information from internal or external sources that is used by the trust algorithm.
Moreover, ZTA relies on Authentication and Identity Management mechanisms to securely verify users and devices. For user authentication, biometric authentication methods [63] such as fingerprint scans, facial recognition, voice recognition, and iris scans provide an additional layer of security by leveraging unique physical characteristics that are difficult to replicate or steal. Advanced methods such as multi-factor authentication (MFA) [61], and identity federation [122], [166] further enhance security while improving user experience. Additionally, innovative techniques such as Physically Unclonable Functions (PUFs), and device fingerprinting are employed for device authentication [61]. PUFs leverages the intrinsic physical properties of hardware to generate unique, tamper-resistant identifiers [81]. Device fingerprinting on the other hand, analyzes hardware and software attributes to create a unique profile for each device [79]. These and other advanced methods ensure that only verified and trustworthy users and devices gain access to critical resources, aligning with ZTA’s zero-trust principles.
@@ 5.3 Enabling Technologies
The realization of ZTA is supported by several enabling technologies that ensure that its principles and requirements can be implemented effectively. These technologies ensure secure, scalable, and resilient operations and management of ZTA in increasingly complex and dynamic environments. Table  4 presents a distribution of the reviewed articles based on these technologies which are critical for ZTA adoption and implementation as discussed in the following subsections. From the table, it is evident that blockchain has been the most used enabling technology by researchers in implementing ZTA frameworks over the years. This is followed by machine learning (ML) and deep learning (DL), and encryption. Blockchain provides immutable and decentralized ledgers which enhances security, efficiency, and trust in digital systems [112], [117] despite challenges such as scalability and regulatory uncertainty. The use of encryption techniques in ZTA is to helps in ensuring the confidentiality of data in all its stages from unauthorized access. On the other hand, ML/DL provides dynamic, context-aware decision-making for threat detection and adaptive access control.
@@ 5.3.2 Behavioral Analytics (BhvA)
Federated Learning is a privacy-preserving machine learning technique that enables the training of models on decentralized datasets so that sensitive information is not compromising [176]. In the context of ZTA, FL is particularly relevant in scenarios where continuous authentication is required while preserving user privacy [134], [30], [69], [171], [68]. For example, the framework proposed in [134] employs FL to authenticate users based on multimodal biometric data. The FL-based approach allows models to be trained on user data which are distributed across various virtual reality (VR) headsets, and hence, eliminates the need for centralized data collection and enhances privacy protection. Similarly, in [30], FL is used to enable distributed training of local models using each device’s unique dataset and submitting the updates to a global model by leveraging Blockchain smart contract, thus, preserving data privacy.
5.3.3 Privacy-preserving Techniques (PrivPres)
@@
Encryption (whose opposite is decryption) in its braod sense involves the conversion of a plaintext to a form that is unreadable or meaningless (called ciphertext) to an unintended recipient using a key [112]. It is used to protect the confidentiality of data at rest (i.e., stored), in motion (i.e., in transmission), or during processing [127], [81]. Encryption can be performed either end-to-end, i.e., the process of encryption and the corresponding decryption are done at the end devices, or point-to-point, which encrypts and decrypts the data at each intermediary device [5]. From the reviewed articles, various studies have proposed the use of different encryption techniques to enhance the robustness of ZTA frameworks in securing access to data from unauthorized entities. For example, in [83], the authors emphasized the use of digital signatures and certificates to enable data privacy where a content publisher uses their private key to digitally sign a user token which ensures that the data (i.e., the published content) is authentic and has not been tempered with. The receiver, on the other hand, uses public key to verify the digital certificates, confirming its issuance by a trusted authority. The ZTA framework in [112] incorporates robust encryption techniques to protect sensitive financial data which establishes a secure environment for transactions while ensuring compliance with industry standards.
Similarly, in [60], the Advanced Encryption Standard (AES) is used to secure manuscript handling between authors, editors, and reviewers. A hash-based key derivation function (HKDF) is used to generate unique keys for each reviewer which ensures that only designated reviewers can access the manuscript, thereby preserving confidentiality during the review process. Also, in [78], a scheme combining zero-trust access control with attribute-based encryption (ABE) to protect against compromised devices in power IoT environments where AES-256 is used to encrypt sensitive data. The Multi-Factor Authentication (MFA) framework designed for zero-trust networks in [107] utilizes asymmetric encryption techniques, such as Elliptic Curve Digital Signature Algorithm (ECDSA) and Elliptic Curve Cryptography (ECC), to ensure the confidentiality of user information throughout the MFA process.
@@ 5.3.4 Decentralized Identity Management Techniques
Identity encompasses a range of characteristics that define a user, such as age, identification, affiliations/organizations to which the user belongs to e.t.c. A framework for federating this identification across several organizations is called identification Federation (IdF) [166]. An authority that authenticates users and provide assertions including the user’s identity and the authority’s signature is called an identity provider or authentication authority [178]. Federated identity management plays a crucial role in enhancing security and user experience within ZTA by streamlining authentication and authorization processes across multiple domains or services. In [166], the authors argue that in a typical ZTA implementations, organizations usually have a single central system that collects contexts from users where as in federated identity, they are distributed across more than one ”Relying Parties (RPs)” and ”Identity Providers (IdPs)”. which can hinder precised authorization decisions. In their proposed ZT Federation framework, they introduced a ”Context Attribute Provider (CAP)” that collects and shares context across the IdF and this enables RPs to leverage a richer set of contextual data for access control.
@@ 5.3.5 Software Defined Networking (SDN)
Another important technology that plays a crucial role in ZTA realization is Software defined Networking (SDN). SDN offers dynamic control over network segments and traffic management which enables precise policy enforcement and supporting the micro-segmentation requirements of ZTA [174], [25]. From the reviewed articles, Usman et al. [174] leverages SDN’s centralized control plane to implement micro-segmentation and dynamically reconfigure traffic flows in IoT environment based on trust levels of nodes. When the trust level of a node falls below a certain threshold, the SDN controller isolates the compromised node by rerouting traffic, and hence, effectively preventing lateral movement of potential attackers. Similarly, in [4], SDN is used to implement micro-segmentation for heterogeneous industrial environments. It enforces security policies at a granular level by isolating network segments and managing encryption processes to contain potential threats. Moreover, in [101], the potential of SDN for establishing a flexible and scalable ZTA in 6G networks is demonstrated. A local SDN controller which is deployed on cloud manages network communities (domains) by collaborating with a third-party security services (TPSSs) to evaluate the trustworthiness of devices requesting access. This consequently mitigates security threats such as malware spread, distributed denial of service (DDoS) attacks, and zero-day exploits.
@@ 5.3.7 ZT Network Access (ZTNA)
ZT Network Access (ZTNA) replaces traditional virtual private networks (VPNs) with identity-aware, granular access to applications and data, aligning with the core tenets of ZTA [125]. While VPNs are widely used to provide secure remote access to organizations’s resources, some security issues are identified which include reliance on traditional PBSM, ”authenticate once, full access to network resources”, and vulnerability to some cyber threat such as phishing attcks, ransomware attacks, malware attacks[181]. Current research focus is towards ZT-enabled solutions. For instance, in [125], an online laboratory access is proposed using ZTNA which addresses the security concerns of VPNs by granting access only to specific applications or resources while adhering to the principle of least privilege. Also, [171] highlighted the significance of ZTNA for enhancing security in cloud environments through continuous authentication, and authorization enforcement, and monitoring of users and devices.
@@ 6.1.1 Fine-grained access control strategies
ZTA is based on the principle of ”never trust, always verify”, where no entity in the organizational network or requesting resources from the organization is assume to be trustworthy until appropriately verified. To effectively establish this principle, it requires security administrators to define and enforce highly specific policies regarding who can access what resources and under what conditions [182]. While fine-grained access control can enhance security by ensuring correct access permissions, it also introduces complexities in management [4]. In other words, defining and maintaining such detailed rules can be resource-intensive, and ensuring that policies are properly enforced at all times requires robust systems and tools. Additionally, the issue can be exacerbated depending on the complexity of the network and the sensitivity of the network resources to be protected. Effective fine-grained access control is more of a post-implementation issue which will increase as the dynamics of the network continue to change.

### v2

@@ 5.1.10 Smart Energy Systems
Smart Energy Systems (SES) integrate modern technologies like the Internet of Things (IoT) [160], cloud computing, and artificial intelligence into existing energy infrastructure. Such systems are meant to maximize energy generation, distribution, and consumption while improving reliability and sustainability of power [161]. However, this interconnectedness brings significant cyber security risks, adversaries could take advantage of vulnerabilities to disrupt services or create economic and societal damage [162]. In essence, ZTA can serve as a powerful model to tackle these security challenges as it guarantees that only authenticated and authorized parties such as devices, users, or systems can access sensitive energy infrastructure [26]. Of the reviewed articles, three specific areas have been identified in which ZTA enhances the security of intelligent energy systems.
@@
Furthermore, Hrishav et al. put forward a ZTA framework based on ABAC to bolster the security of AMI within smart grids [167]. The framework employs a trust score-based methodology, where each attribute (a set of 15 compiled attributes) receives a trust score which signifies the level of trust put in it and a corresponding threshold score which is represents the minimum required trust score for a given attribute to be taken as reliable. The authors consider a case study using five attributes which include ”Smart Meter ID, Device Version, Location, Challenge Response Pairs (CRPs)”, and ”Network Used” where a smart meter starts an access request to transmit usage data to a power supply company. Network access is granted only when the total trust score value is more than the total threshold value of the attributes. The case study shows the importance of each attribute in the access verification process within the AMI system. Also, the verification mechanism, driven by trust scores, ensures secure and reliable communication between smart meters and the utility company which effectively prevents unauthorized access and maintaining data integrity.
@@
Power IoT devices such as sensors and controllers are used in smart grids for optimal energy management, however, these devices are also vulnerable to cyber threats like malware and denial of service attacks. Through microsegmentation and other strategies, ZTA can protect power IoT systems such that a compromised device cannot propagate its effect to the full network [125]. Based on the key principles of ZT and the features of the power IoT, the authors in [160] proposed the ZT framework that can evaluate the security state of power IoT derived from the behaviuor and characteristics of the terminals. Through a comparative study based on simulation results and stochastic Petri nets, they prove that the proposed framework can be used to effectively solve network security problems, and significantly improve the security protection of the power IoT. Similarly, to fortify the access control and protect the power IoT environment against and disgruntle users and compromised devices, a ZTA framework based on attribute-based encryption method is proposed in [84]. It is made up of three interconnected modules - ”control plane” acting as the intelligent core of the framework and responsible for making real-time access decisions through dynamic trust evaluation and security access control components, ”data plane” which receives and executes the access control decisions from the control plane, and the ”identity security infrastructure” which handles the registration and authentication process of users and devices, manages their attributes, and defines the access policies that govern access to specific resources. The interaction of these three modules ensures that only trusted entities with validated identities and acceptable trust scores are granted access to sensitive data within the power IoT environment. Experimental results show that our framework is effective against the identified security vulnerabilities.
@@
Development in information technology (IT) has both significantly improved the power grids and systems operations and exposed the power infrastructure to diverse security challenges which the traditional security controls cannot be able to cope with effectively [168]. Hence, robust and dynamic security solutions such as ZTA frameworks are needed to secure data and information and terminal devices in the power system networks [169]. For example, in [169], the authors proposed a three-pronged approach to developing a ZTA framework for securing terminal access in distributed energy grids: ”Security modules”, like ”Trusted Platform Modules (TPMs) or Hardware Security Modules (HSMs)”, act as trust anchors for the system storing important information like digital certificates or terminal keys. ”Identity authentication” which relies on lightweight one-time digital signatures, ensuring real-time authentication and enhancing security by generating a new signature for each request. ”Security detection” employs data stream classification to continuously monitor the security of terminals, facilitating real-time analysis and anomaly detection. Through comparative analysis, the proposed framework is shown to outperformed traditional methods and can address the security challenges of decentralized power grids.
@@ 5.2 Application Requirements
In order to successfully deploy ZTA, specific baseline requirements must be met. Microsegmentation is a cornerstone, dividing networks into smaller, isolated zones to minimize the impact of breaches and prevent the lateral movement of threats [14], [112]. This includes specifying trust dimensions for each access request such that access is granted only when defined security parameters are satisfied [21]. Continuous monitoring and verification is essential to the ZTA model by analyzing user and device activities in real time to detect anomalies and enforce security policies dynamically [7]. This process relies on monitoring and analytics solutions, which involve the use of Security Information and Event Management (SIEM) [79] and Security Orchestration, Automation and Response (SOAR) systems [170] to provide centralized log data aggregation from systems and appliances, real-time threat detection, and automated response [171]. Security concepts such as threat intelligence platforms and activity log management further enhance situational awareness by providing insights into emerging threats and enabling traceability for incident response [6]. This includes information from internal or external sources that is used by the trust algorithm.
Moreover, ZTA relies on Authentication and Identity Management mechanisms to securely verify users and devices. For user authentication, biometric authentication methods [69] such as fingerprint scans, facial recognition, voice recognition, and iris scans provide an additional layer of security by leveraging unique physical characteristics that are difficult to replicate or steal. Advanced methods such as multifactor authentication (MFA) [67], and identity federation [128], [172] further enhance security while improving user experience. Additionally, innovative techniques such as Physically Unclonable Functions (PUFs), and device fingerprinting are employed for device authentication [67]. PUFs leverage the intrinsic physical properties of hardware to generate unique, tamper-resistant identifiers [87]. Device fingerprinting on the other hand, analyzes hardware and software attributes to create a unique profile for each device [85]. These and other advanced methods ensure that only verified and trustworthy users and devices gain access to critical resources, aligning with ZTA’s zero-trust principles.
@@ 5.3 Enabling Technologies
The realization of ZTA is supported by several enabling technologies that ensure that its principles and requirements can be implemented effectively. These technologies ensure secure, scalable, and resilient operations and management of ZTA in increasingly complex and dynamic environments. Table  4 presents a distribution of the reviewed articles based on these technologies which are critical for ZTA adoption and implementation as discussed in the following subsections. From the table, it is evident that blockchain has been the most used enabling technology by researchers in implementing ZTA frameworks over the years. This is followed by machine learning (ML) and deep learning (DL), and encryption. Blockchain provides immutable and decentralized ledgers which enhances security, efficiency, and trust in digital systems [118], [123] despite challenges such as scalability and regulatory uncertainty. The use of encryption techniques in ZTA is to help in ensuring the confidentiality of data in all its stages from unauthorized access. On the other hand, ML/DL provides dynamic, context-aware decision-making for threat detection and adaptive access control.
@@ 5.3.2 Behavioral Analytics (BhvA)
Federated Learning is a privacy-preserving machine learning technique that enables the training of models on decentralized datasets so that sensitive information is not compromising [182]. In the context of ZTA, FL is particularly relevant in scenarios where continuous authentication is required while preserving user privacy [140], [30], [75], [177], [74]. For example, the framework proposed in [140] employs FL to authenticate users based on multimodal biometric data. The FL-based approach allows models to be trained on user data which are distributed across various virtual reality (VR) headsets, and hence eliminates the need for centralized data collection and enhances privacy protection. Similarly, in [30], FL is used to enable distributed training of local models using each device’s unique dataset and submitting the updates to a global model by leveraging Blockchain smart contract, thus preserving data privacy.
5.3.3 Privacy-preserving Techniques
@@
Encryption (whose opposite is decryption) in its braod sense involves the conversion of a plaintext to a form that is unreadable or meaningless (called ciphertext) to an unintended recipient using a key [118]. It is used to protect the confidentiality of data at rest (i.e., stored), in motion (i.e., in transmission), or during processing [133], [87]. Encryption can be performed either end-to-end, i.e., the process of encryption and the corresponding decryption are done at the end devices, or point-to-point, which encrypts and decrypts the data at each intermediary device [5]. From the reviewed articles, various studies have proposed the use of different encryption techniques to enhance the robustness of ZTA frameworks in securing access to data from unauthorized entities. For example, in [89], the authors emphasized the use of digital signatures and certificates to enable data privacy where a content publisher uses their private key to digitally sign a user token which ensures that the data (that is, the published content) is authentic and has not been tempered with. The receiver, on the other hand, uses public key to verify the digital certificates, confirming its issuance by a trusted authority. The ZTA framework in [118] incorporates robust encryption techniques to protect sensitive financial data which establish a secure environment for transactions while ensuring compliance with industry standards.
Similarly, in [66], the Advanced Encryption Standard (AES) is used to secure manuscript handling between authors, editors, and reviewers. A hash-based key derivation function (HKDF) is used to generate unique keys for each reviewer which ensures that only designated reviewers can access the manuscript, thereby preserving confidentiality during the review process. Also, in [84], a scheme combining zero-trust access control with attribute-based encryption (ABE) to protect against compromised devices in power IoT environments where AES-256 is used to encrypt sensitive data. The multifactor Authentication (MFA) framework designed for zero-trust networks in [113] utilizes asymmetric encryption techniques, such as Elliptic Curve Digital Signature Algorithm (ECDSA) and Elliptic Curve Cryptography (ECC), to ensure the confidentiality of user information throughout the MFA process.
@@ 5.3.4 Decentralized Identity Management Techniques
Identity encompasses a range of characteristics that define a user, such as age, identification, affiliations/organizations to which the user belongs to etc. A framework for federating this identification across several organizations is called identification Federation (IdF) [172]. An authority that authenticates users and provide assertions including the user’s identity and the authority’s signature is called an identity provider or authentication authority [184]. Federated identity management plays a crucial role in enhancing security and user experience within ZTA by streamlining authentication and authorization processes across multiple domains or services. In [172], the authors argue that in a typical ZTA implementations, organizations usually have a single central system that collects contexts from users where as in federated identity, they are distributed across more than one ”Relying Parties (RPs)” and ”Identity Providers (IdPs)”. which can hinder precised authorization decisions. In their proposed ZT Federation framework, they introduced a ”Context Attribute Provider (CAP)” that collects and shares context across the IdF and this enables RPs to leverage a richer set of contextual data for access control.
@@ 5.3.5 Software Defined Networking (SDN)
Another important technology that plays a crucial role in ZTA realization is Software defined Networking (SDN). SDN offers dynamic control over network segments and traffic management which enables precise policy enforcement and supporting the micro-segmentation requirements of ZTA [180], [25]. From the reviewed articles, Usman et al. [180] leverages SDN’s centralized control plane to implement micro-segmentation and dynamically reconfigure traffic flows in IoT environment based on trust levels of nodes. When the trust level of a node falls below a certain threshold, the SDN controller isolates the compromised node by rerouting traffic, and hence, effectively preventing lateral movement of potential attackers. Similarly, in [4], SDN is used to implement microsegmentation for heterogeneous industrial environments. It enforces security policies at a granular level by isolating network segments and managing encryption processes to contain potential threats. Moreover, in [107], the potential of SDN for establishing a flexible and scalable ZTA in 6G networks is demonstrated. A local SDN controller which is deployed on cloud manages network communities (domains) by collaborating with a third-party security services (TPSSs) to evaluate the trustworthiness of devices requesting access. This consequently mitigates security threats such as malware spread, distributed denial of service (DDoS) attacks, and zero-day exploits.
@@ 5.3.7 ZT Network Access (ZTNA)
ZT Network Access (ZTNA) replaces traditional virtual private networks (VPNs) with identity-aware, granular access to applications and data, aligning with the core tenets of ZTA [131]. While VPNs are widely used to provide secure remote access to organizations’ resources, some security issues are identified which include reliance on traditional PBSM, ”authenticate once, full access to network resources”, and vulnerability to some cyber threat such as phishing attcks, ransomware attacks, malware attacks[187]. Current research focus is towards ZT-enabled solutions. For instance, in [131], an online laboratory access is proposed using ZTNA which addresses the security concerns of VPNs by granting access only to specific applications or resources while adhering to the principle of least privilege. Also, [177] highlighted the significance of ZTNA for enhancing security in cloud environments through continuous authentication, and authorization enforcement, and monitoring of users and devices.
@@ 6.1.1 Fine-grained access control strategies
ZTA is based on the principle of ”never trust, always verify”, where no entity in the organizational network or requesting resources from the organization is assumed to be trustworthy until appropriately verified. To effectively establish this principle, it requires security administrators to define and enforce highly specific policies regarding who can access what resources and under what conditions [188]. While fine-grained access control can enhance security by ensuring correct access permissions, it also introduces complexities in management [4]. In other words, defining and maintaining such detailed rules can be resource-intensive, and ensuring that policies are properly enforced at all times requires robust systems and tools. Additionally, the issue can be exacerbated depending on the complexity of the network and the sensitivity of the network resources to be protected. Effective fine-grained access control is more of a post-implementation issue which will increase as the dynamics of the network continues to change.